OpenStudy (anonymous):

HELP!!!!!!!!!! Solve the following system. x − y = 10 x + y = 4 (−3, 7) (−7, 3) (7, 3) (7, −3)

OpenStudy (anonymous):

(7,-3)

OpenStudy (anonymous):

can you show your work please thnxs

OpenStudy (anonymous):

x=10+y 10+y+y=4 => 10-4=-2y=> 6=-2y=> y=-3

OpenStudy (anonymous):

x=10-3=7

OpenStudy (anonymous):

thank you